OpenStudy (anonymous):

81w^6-3p^6 factor medals given

OpenStudy (anonymous):

lol not again...this is based on the same formula..why not try it urself ?

OpenStudy (anonymous):

I have tried. I just can't get the right answer. All the ones I have posted I have done several times on my own.

OpenStudy (anonymous):

take the 3 out

OpenStudy (anonymous):

look for common factors first, you finish it

myininaya (myininaya):

3(27w^6-p^6)

OpenStudy (anonymous):

this is why you should be familiar with the squares and cubes of the numbers 1-12

OpenStudy (anonymous):

so its easy to find patterns.

myininaya (myininaya):

or \[(9w^3-\sqrt{3}p^3)(9w^3+\sqrt{3}p^3)\]

OpenStudy (anonymous):

3(3w^2-p^2)(9w^4+3w^2*p^2+p^4) thats it

myininaya (myininaya):

\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

even then that first difference of sqaures should be factored further

myininaya (myininaya):

you also have a different of cubes formula and a sum of cubes formula

OpenStudy (anonymous):

3w^2 -p^2 , shouldnt leave as this, everything should be linear or irreduciable quadratic.

OpenStudy (anonymous):

whats the problem with my answer then? \[a ^{3}-b ^{3}=(a ^{2}+ab+b ^{2}) (a-b)\]

OpenStudy (anonymous):

thanks guys. all these formulas! I can't keep up.